what is the equation of the line passing through the point (1, -5) and perpendicular to y = 1/5x - 10.
Answer by pablo9619(2)

(1,-5) perpendicular to y=1/5x-10
perpendicular lines have slopes that are opposite
1/5 -5/1
y=mx+b
-5=-5(1)+b
-5=-5+b
+5 +5
0=b
y=-5x+0